If x^2 + x – 6 = 0, then x is:
Answer by texttutoring(324)

x^2 + x – 6 = 0

This is a quadratic (there is an x^2 term), so you need to solve it by factoring. You will get two answers.

You need to find two numbers that multiply to -6 and also add up to +1.
Those numbers are 3 and -2.

x^2 + x – 6 = 0
(x+3)(x-2)=0

So, either (x+3)=0 or (x-2)=0.

If x+3=0, x=-3

If x-2=0, x=2

x=-3, x=2 are your answers.